OpenStudy (anonymous):

f(x)=4⋅x+2+2h(x+1) h(n)=3⋅n h(f(-2))=?

OpenStudy (asadkarim7):

1st take f(-2) so = 4. -2+2+2h(-2+1) = -8 + 2+2h (-1) = -6 -2h it means h ( -6-2h )= 3 . -6-2h -18 -2h

OpenStudy (anonymous):

1. you calculate 2h(x+1)=2*3(x+1)=6x+6 2. f(-2)=-8+2-12+6=-12 3. h(-12)=-36